Suppose a 10-year, 10 percent, semiannual coupon bond with a par value of R1 000
is currently selling for R1 135.90, producing a nominal yield to maturity of 8 percent.
However, the bond can be called after 5 years for a price of R1 050.
i) (3 points) What is the bond’s nominal yield to call?
ii) (2 points) If you bought this bond, do you think you would be more likely to earn the
YTM or the YTC? Why?
1
Expert's answer
2020-09-06T13:31:43-0400

i)"YTC=\frac{C+\frac{(FP-MP)}{n}}{\frac{FP+MP}{2}}"

C=50

FP=1050

MP=1135.9

n=10

"YTC=\frac{50+\frac{(1050-1135.9)}{n}}{\frac{1050+1135.9}{2}}=0.037" or 3.7%

"0.037\times2=7.5"

ii)YTC=7.5

YTM=8


coupon rate 10% and YTC=rd=7.5%

if coupon rate >rd bond sells at a premium.

It also depends on fluctuations in interest rates from today to the maturity date, however, you would expect to earn on YTC on premium bonds, as well as earn ytmon par and discount bonds